OpenStudy (anonymous):

Medals and fan rewarded! Find all solutions in the interval [0, 2π). 2 sin2x = sin x

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

2 sin^2x = sin x****

OpenStudy (anonymous):

@bina_love do you know hot to solve this?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

subtract sin x from both sides and factor it out from the left

OpenStudy (anonymous):

These are my answer choices A. x = pi/3 , 2pi/3 B. x = pi/2 , 3pi/2 , pi/3 , 2pi/3 C. x = 0 , pi , pi/6 , 5pi/6 D. x = pi/6 , 5pi/6

OpenStudy (anonymous):

I dont see how i can get any of those answers from doing what you told me to do? @hartnn

hartnn (hartnn):

i'll tell you how, firstly could you do what i said ?

OpenStudy (anonymous):

I would be left with 2 sin x correct?

hartnn (hartnn):

what did u get after subtracting sin x from bot hsides ?

OpenStudy (anonymous):

2 sin x?

hartnn (hartnn):

noooo... 2sin^2 x = sin x so, 2 sin^2 x - sin x = 0 now factor out sin x from here

OpenStudy (anonymous):

i should end up with x ( 2x -1 ) = 0 correct?

hartnn (hartnn):

yeah, actually sin x (2 sin x-1) = 0 thats the factored form

hartnn (hartnn):

so either sin x = 0 or 2 sin x -1 = 0 got this ?

OpenStudy (anonymous):

Yeah okay now what

hartnn (hartnn):

when will sin x be = 0 ?

OpenStudy (anonymous):

Im not sure, can you just show me how to do it step by step and then ill ask you uestions afterwards? This is going to take forever, its just a practice problem regardless.

OpenStudy (anonymous):

Questions*

hartnn (hartnn):

sin x will be 0 when x=0 !! or x= pi, 2pi ...

OpenStudy (anonymous):

But thats not one of my answer choices?

hartnn (hartnn):

in 0 to 2pi we will have 0 and pi now, for 2sin x -1=0 sin x =1/2 , right ?? for which angles will sin x be = 1/2 ??

OpenStudy (anonymous):

x?

hartnn (hartnn):

i asked for which angles ? ok, sin x is 1/2 for x = pi/6 , 5pi/6 so our 4 angles will be x = 0 , pi , pi/6 , 5pi/6

OpenStudy (anonymous):

Thank you so much